A chemist has 1000 g of a solution that is 40% acid. How many grams of water must be added to reduce the acidity to 25%?
Answer by checkley71(8403)

1000*.4=x+.25(1000+x)
400=x+250+.25x
1.25x=150
x=150/1.25
x=120 g of pure water needs to be added.
proof
400=120+250+30
400=400